Three people share $5992 so that the first gets twice as much as the second, and the second gets twice as much as the third. How much does the first get?
Answer by ankor@dixie-net.com(22740)

Three people share $5992 so that the first gets twice as much as the second, and the second gets twice as much as the third.
How much does the first get?
:
let the 3 amts be a, b, c
" the first gets twice as much as the second,"
a = 2b
"the second gets twice as much as the third."
b = 2c
divide both sides by 2
c = .5b
:
a + b + c = 5992
replace a and c
2b + b + .5b = 5992
3.5b = 5992
b = 5992/3.5
b = 1712
:
How much does the first get?
a = 2(1712)
a = $3424